Retractable Technologies just paid a dividend of D_0 = $2.20 and sells for $40 per share. If the company has a constant growth rate of 6% per year, what is their total expected return?
Answer #1

Total expected return=Dividend payment in the coming year/Current stock price + Constant growth rate
Dividend payment in the coming year= Dividend paid this year*(1+ growth rate)
Given that the dividend paid this year=$2.2 and the constant growth rate=6%

Dividend payment in the next year=$2.2*(1+6%)=$2.2*(1.06)=$2.332
Current stock price=$40 per share
Total expected return=$2.332/$40+6%=0.0583+0.06=0.1183 or 11.83%